OpenStudy (anonymous):

Solve. 9x = 3x – 36 A. –6 B. –4 C. 4 D. 6

jigglypuff314 (jigglypuff314):

9x = 3x – 36 subtract 3x from both sides -3x -3x 6x = -36 divide both sides by 6 /6 /6 x = - 36/6 =

OpenStudy (anonymous):



OpenStudy (anonymous):